OpenStudy (anonymous):

in an election, the winning candidate reciived 56,000 votes which represented 70% of the electrotates. The only other candidate secured 15% votes of the electrotates calculate: 1.. the number of the electorates 2.. the number of people who did not cast their votes.

OpenStudy (anonymous):

Give me a good answer and i'll help :)

OpenStudy (anonymous):

the answer is (80000 , 12000)

OpenStudy (anonymous):

I mean click my good answer button so i level up

OpenStudy (anonymous):

lol, what is the problem Ra Abi?

OpenStudy (anonymous):

i want steps for ths questions and Dangolbery how do i do that?

OpenStudy (anonymous):

doneDangolbery

OpenStudy (anonymous):

56000 is 70% of the votes 56000 / 70 gives what 1% is equal too 56000/70 * 100 gives the total votes that can be cast

OpenStudy (anonymous):

A 70 % B 15 % C = 100 - 70 - 15 C = 15 56000/70 * 15

OpenStudy (anonymous):

56, 000 = 0.7 (70%) 56,000/0.7 = 80,000 total people in the electorate Other candidate got 0.15*80,000 = 12,000 votes Total people who voted is 56,000 (winning candidate) 12,000 (other candidate) 80,000 - 56,000 - 12,000 = 12,000 remaining people who did not vote.

OpenStudy (anonymous):

thank u sooo much:)

OpenStudy (anonymous):

You're welcome :)